Identify the conic section, write each in standard form, and graph:
4x^2-9y^2+y+3=0

Answer by stanbon(48581)

Identify the conic section, write each in standard form, and graph:
4x^2-9y^2+y+3=0
Complete the square on the x-terms and on the y-terms:
4(x^2) -9(y^2-(1/9)y = -3
4(x^2) -9(y^2-(1/9)y+(1/18)^2) = -3-9*(1/18)^2
4(x^2) -9(y-(1/18))^2 = -109/36
Divide both sides by -109/36 to get:
-(x-0)^2]/(109/144) +(y-(1/18))^2/(109/324) = 1
------------------
It's a hyperbola.
Can't be graphed on this site as it is not a function.
==================
Cheers,
Stan H.